OpenStudy (anonymous):

SOLVE the equation (3x-1)/(x+7) -(1)/(3x+2) = (5-x)/(3x^2+23x+14)

OpenStudy (anonymous):

set (3x^2+23x+14)=(7+x) (2+3x)

OpenStudy (anonymous):

why would I do that? When I first saw the problem, I thought that I should clear fractions with a common LCD The common LCD is (3x+2)(x+7) so multiply everything by that. then I got (3x-1)(3x+2)-(x+7)=(5-x) when I foil and combine like terms, setting the entire thing equal to 0, I got: (3x+5)(3x-2)=0 so x= 2/3 and -5/3 That is what I have...not sure if it is correct tho..